# Phonon-Coupled Hole-Spin Qubits in High-Purity Germanium: Design and Modeling of a Scalable Architecture

### v1

@@ I Introduction
To place this platform within the broader quantum device landscape, Fig. 1 compares various qubit modalities in terms of coherence time, near-term scalability, and gate speed. Ge hole-spin qubits with phonon coupling uniquely occupy a favorable region combining long coherence with fast, purely electrical manipulation. Compared to superconducting qubits and trapped ions, the phonon-enhanced Ge platform offers a desirable balance that directly supports scalable quantum processor design.
Figure 1: Visual comparison of leading qubit modalities based on coherence time, gate speed, and near-term scalability. Note that the comparison reflects current experimental capabilities [1, 2, 8, 9], not long-term potential. Ge-based hole-spin qubits with phonon coupling occupy a favorable region, underscoring their potential for scalable and high-fidelity quantum processing.
@@ II Spin-Phonon Coupling in Germanium Quantum Dots
Figure 2: Schematic showing spin-phonon coupling in a Ge quantum dot. Under an applied magnetic field along [001], spin states split by the Zeeman energy $\Delta E$, and phonons with matching energy $E_{ph}=\hbar\omega$ enable spin-flip transitions. These transitions are mediated via spin-orbit coupling and enhanced by heavy-hole/light-hole mixing in low-symmetry environments.
@@ III Phonon-Assisted $T_{1}$ Relaxation
In addition to enabling all-electrical spin control, spin-phonon coupling plays a key role in limiting qubit performance by contributing to spin relaxation. Specifically, the longitudinal relaxation time $T_{1}$, which characterizes the decay of an excited spin state to its ground state, is predominantly governed by phonon-mediated processes in systems with strong spin-orbit coupling (SOC). Understanding and engineering this relaxation mechanism is critical for achieving long-lived, high-fidelity qubits in Ge quantum dot platforms.
@@
$\frac{1}{T_{1}}\propto|M_{\text{sp}}|^{2}D(\omega),$ (2)
@@ IV Phononic Crystal Cavity for Control and Coupling
The interaction between the qubit and cavity mode is enhanced by strong spin-orbit coupling in the valence band of Ge, and by spatial overlap between the phonon mode and the confined hole wavefunction. Electric field tuning allows real-time modulation of the g-factor [7], enabling dynamic control of the qubit–cavity detuning. This feature is essential for implementing phonon-mediated two-qubit gates, dispersive readout schemes, and coherent state transfer between distant nodes [37, 32, 38].
@@
To enable phonon-mediated control of hole-spin qubits, we have designed a two-dimensional (2-D) PnC based on Ge that supports an absolute bandgap centered in the 4–6 GHz range, with a width of approximately 2–3 GHz. This frequency range corresponds to thermal phonon energies near 0.3 K and aligns with the operating environment of dilution refrigerators, which typically operate at temperatures around 100 mK. Within this regime, thermal phonons can contribute to decoherence, and thus, phononic bandgap structures act as spectral filters that suppress unwanted phonon modes while enabling selective transmission of engineered phononic states. This filtering capability enhances coherence and allows for precise, long-range phonon-mediated qubit control in Ge-based quantum architectures.
@@
In this work, we estimate the relative strength of spin–phonon coupling $g_{\text{sp}}$ by evaluating the spatial overlap between the phononic strain field and the qubit’s envelope wavefunction. Using a Gaussian approximation for the hole-spin qubit and the simulated strain profile of a 6 GHz confined phonon mode, we compute a normalized overlap integral of approximately $2.6\times 10^{-2}$. This value is over four orders of magnitude greater than that obtained at 20 GHz, reflecting the fact that lower-frequency phonons, which exhibit broader and smoother spatial profiles, are better matched to the spatial extent of gate-defined quantum dots. The improved spatial overlap reduces destructive interference within the qubit region and enables stronger, more coherent coupling between the spin and the phonon field.
@@ V Experimental Considerations
The experimental realization of phonon-coupled hole-spin qubits in Ge requires careful coordination of high-purity material growth, precision nanofabrication, and cryogenic quantum measurements. The foundation of this platform lies in the exceptional material quality available at USD, where detector-grade, high-purity Ge crystals are synthesized in-house. These crystals exhibit impurity concentrations below $10^{10}~{}\text{cm}^{-3}$, substantially suppressing charge noise and impurity scattering—key sources of qubit decoherence [13, 14, 15, 16, 17, 18, 19]. This vertically integrated purification and crystal growth capability uniquely positions USD to fabricate ultrapure substrates tailored for quantum devices.
@@
Quantum dots are electrostatically defined using overlapping gate architectures fabricated by high-resolution electron-beam lithography. A high-$\kappa$ dielectric layer (e.g., HfO2) deposited via atomic layer deposition isolates the gates from the Ge surface. The gate stack typically includes accumulation gates to form the quantum channel, barrier gates to define tunnel junctions, and plunger gates to tune the dot occupancy. Vertical electric fields generated by the gates modulate the hole g-factor through Rashba spin-orbit interaction, enabling electric-dipole spin resonance (EDSR) without requiring oscillating magnetic fields [7, 8].
@@ VI Device Design Concept
Figure 10: Schematic cross-sectional diagram of a Ge quantum dot qubit platform. A single hole is confined within a gate-defined Ge quantum dot embedded in a strained Ge layer, grown atop a relaxed Si1–xGex virtual substrate. The structure includes a cap layer and top accumulation gate to provide vertical confinement and electrostatic control. The epitaxial strain enhances heavy-hole character and spin–orbit coupling, enabling fast, electrically tunable qubit operation. The coordinate axes indicate the crystallographic directions relevant to quantum confinement and gating.
@@
Figure 11: Schematic illustration of a hole-spin qubit confined in a strained Ge quantum dot embedded within a Ge quantum well. Gate electrodes define the quantum dot electrostatically, while a nearby PnCC enables phonon coupling for coherent manipulation. The heterostructure is grown on a relaxed Si1-xGex buffer layer to induce strain in the quantum well, enhancing the confinement and coherence properties of the hole spin.
@@ VII.2 Quantum Dot Behavior and Spin-Orbit Coupling
where $g_{\text{eff}}(E_{z})$ is the electric-field-dependent effective g-factor, $\mu_{B}$ is the Bohr magneton, and $B$ is the applied magnetic field. The spin–phonon coupling strength $g_{sp}$ is computed as [35, 41]:
$g_{sp}\propto\left|\langle\psi_{\uparrow}(\vec{r})\,|\,H_{\text{so}}\,|\,\psi_{\downarrow}(\vec{r})\rangle\right|,$ (11)
where $H_{\text{so}}$ denotes the strain-modulated spin–orbit Hamiltonian and $\psi_{\uparrow,\downarrow}(\vec{r})$ are the spin-resolved qubit eigenstates. This formulation is consistent with the spatial overlap expression in Eq. (3), providing a unified framework for describing spin–phonon interactions in Ge quantum dots.
Building on a multiband $\mathbf{k}\cdot\mathbf{p}$ framework and self-consistent Poisson–Schrödinger solvers [20, 22], the simulation framework described above was employed to model the dependence of $g_{\text{eff}}$ and $g_{sp}$ on the vertical electric field. Figure 12 shows that increasing $E_{z}$ reduces $g_{\text{eff}}$ from 2.0 to 1.3 and simultaneously enhances $g_{sp}$ from 0.5 to over 6 MHz. This tunability is critical for fast EDSR, as it allows dynamic optimization of qubit control fidelity and gate speed. Quantitative simulation results are summarized in Table 3, confirming the trends observed in the figure.
@@
Table 3: Simulated values of effective g-factor and phonon coupling strength (MHz) as a function of electric field.
E-Field (MV/m) g-factor Coupling Strength (MHz)
0.0 2.00 0.5
0.2 1.86 1.2
0.4 1.72 2.1
0.6 1.58 3.5
0.8 1.44 5.0
1.0 1.30 6.3
@@ VII.3 Phonon Dynamics: Relaxation and Confinement
We employed the previously described simulation framework to model spin relaxation dynamics and phonon confinement. The longitudinal relaxation time $T_{1}$ is governed by Fermi’s golden rule [10, 11]:
$\frac{1}{T_{1}}=\frac{2\pi}{\hbar}|M_{sp}|^{2}D(\omega),$ (12)
where $M_{sp}$ is the spin-phonon matrix element and $D(\omega)$ the phonon density of states. In the low-temperature acoustic limit, $D(\omega)\sim\omega^{2}$, leading to shorter $T_{1}$ at higher phonon frequencies [10, 11]. Note that $M_{sp}$ in Eq. (12) and Eq. (2) corresponds to the spin–phonon matrix element derived from $H_{\text{so}}(u(\vec{r}))$, as defined in Eqs. (11) and (3).
@@
where $\omega$ is the mode frequency and $\Delta\omega$ the linewidth obtained from eigenmode simulations. As shown in Figure 13, maximum coherence and cavity performance occur near 6 GHz, with simulated $T_{1}\approx 1$ ms and $Q\approx 18000$. The numerical values corresponding to these trends are listed in Table 3, providing quantitative validation of the simulation results shown in the figure.
@@
Table 4: Simulated $T_{1}$ relaxation and Q-factors as a function of phonon frequency.
Phonon Frequency (GHz) T1 (ms) Q-Factor
2 9.00 10000
3 4.00 13500
4 2.25 16000
5 1.44 17500
6 1.00 18000
7 0.74 17500
8 0.56 16000

### v2

@@ I Introduction
To contextualize the proposed Ge-based phonon-coupled hole-spin qubit within the broader quantum computing landscape, Fig. 1 presents a comparative overview of several leading qubit modalities, using coherence time and gate speed as primary performance benchmarks. The figure highlights where current experimental realizations stand with respect to these two key metrics.
Figure 1: Comparison of representative qubit technologies based on coherence time and gate speed, reflecting current experimental capabilities [1, 2, 8, 9]. Ge-based hole-spin qubits with phonon coupling are positioned in a favorable regime that combines fast, all-electrical control with long coherence, underscoring their promise for high-fidelity quantum information processing.
While scalability is not depicted in Fig. 1, it is an equally important consideration in evaluating qubit platforms. In this work, scalability is treated as a qualitative metric that reflects the practical feasibility of large-scale quantum processor integration. This includes factors such as compatibility with planar CMOS fabrication, support for densely packed qubit arrays, use of purely electrical control mechanisms, and architectural suitability for parallel operations and quantum error correction. The discussion in the main text draws upon established criteria, particularly those outlined by Scappucci et al. [9], to assess the integration-readiness of various platforms.
Among the technologies compared, Ge-based hole-spin qubits with phonon coupling offer a unique combination of millisecond-scale coherence times and nanosecond-scale gate speeds, a regime that enables fast, coherent control while maintaining compatibility with scalable semiconductor manufacturing. In contrast to platforms such as trapped ions or superconducting qubits, which face inherent constraints from optical systems or complex cryogenic wiring, the proposed Ge platform leverages phonon-engineered spin-orbit interactions and mature fabrication techniques to simultaneously optimize coherence, control speed, and integration potential.
@@ II Spin-Phonon Coupling in Germanium Quantum Dots
Figure 2: Schematic showing spin-phonon coupling in a Ge quantum dot. Under an applied magnetic field along [001], spin states split by the Zeeman energy $\Delta E$, and phonons with matching energy $E_{ph}=\hbar\omega$ enable spin-flip transitions. These transitions are mediated via spin-orbit coupling and enhanced by heavy-hole/light-hole mixing in low-symmetry environments. Here, $\lvert v\rangle$ and $\lvert w\rangle$ denote the spin-down and spin-up eigenstates of the spin qubit under an external magnetic field, corresponding to the spin-split states in the Ge quantum dot.
While Eq. (1) assumes an isotropic $g$-factor for simplicity, we note that in strained Ge quantum wells, the $g$-factor and spin–phonon coupling are in fact anisotropic [10, 23]. This anisotropy arises from the complex valence band structure and spin–orbit interactions, which are sensitive to the orientation of the applied magnetic and electric fields relative to the crystallographic axes. Specifically, both the magnitude and sign of the effective $g$-factor can vary with the polar and azimuthal angles of the field vectors [24]. As such, optimal spin–phonon coupling is generally achieved when the magnetic field is aligned along the growth direction (typically [001]) and the electric field lies in the quantum well plane. Misalignment can lead to reduced coupling efficiency or modified spin splitting, which may affect gate fidelity and readout. In future work, the angular dependence of $g_{\mathrm{sp}}$ and its implications for qubit control fidelity will be modeled more explicitly.
@@ III Phonon-Assisted $T_{1}$ Relaxation
In addition to enabling all-electrical spin control, spin-phonon coupling plays a key role in limiting qubit performance by contributing to spin relaxation. Specifically, the longitudinal relaxation time $T_{1}$, which characterizes the decay of an excited spin state to its ground state, is predominantly governed by phonon-mediated processes in systems with strong SOC. Understanding and engineering this relaxation mechanism is critical for achieving long-lived, high-fidelity qubits in Ge quantum dot platforms.
@@
$\frac{1}{T_{1}}=\frac{2\pi}{\hbar}|M_{\text{sp}}|^{2}D(\omega),$ (2)
@@ IV Phononic Crystal Cavity for Control and Coupling
The interaction between the qubit and the cavity mode is strengthened by the strong spin–orbit coupling intrinsic to Ge’s valence band, along with the spatial overlap between the phonon mode and the confined hole wavefunction. Real-time modulation of the qubit–cavity detuning is achieved through electric field tuning of the $g$-factor [7], providing dynamic control over qubit–phonon interactions. This tunability is crucial for enabling phonon-mediated two-qubit gates, dispersive readout schemes, and coherent state transfer across distant nodes [39, 34, 40].
Building on this foundation, we describe how the PnCC architecture facilitates coherent coupling between multiple Ge-based hole-spin qubits to support two-qubit gate operations. For single-qubit control, the platform employs electric dipole spin resonance (EDSR), which harnesses the strong spin–orbit interaction in strained Ge quantum wells. Oscillating gate voltages drive spin rotations, eliminating the need for microwave magnetic fields [7, 8]. Simulations and prior experimental results demonstrate gate times in the 10–50 ns range, enabling fast, all-electrical qubit manipulation well within the available coherence time.
For two-qubit operations, we consider a phonon-mediated dispersive coupling mechanism wherein two spatially separated hole-spin qubits interact through a shared confined phonon mode localized in the PnCC. This mode serves as a virtual mediator analogous to a cavity photon in circuit QED. When both qubits are coupled to the same phonon mode with coupling strength $g_{\text{sp}}$ and are operated at a large detuning $\Delta$ from the phonon resonance, the resulting second-order perturbative interaction yields the effective Hamiltonian:
$H_{\text{eff}}=\frac{g_{\text{sp}}^{2}}{\Delta}\left(\sigma_{+}^{(1)}\sigma_{-}^{(2)}+\sigma_{-}^{(1)}\sigma_{+}^{(2)}\right),$ (3)
where $\sigma_{+}^{(i)}=|1\rangle_{i}\langle 0|$ and $\sigma_{-}^{(i)}=|0\rangle_{i}\langle 1|$ are the spin raising and lowering operators for qubit $i$.
This Hamiltonian represents an effective XY-type interaction, commonly encountered in cavity and circuit QED systems [41, 42]. Although it is not a direct exchange interaction—which requires significant overlap of wavefunctions—it produces a functionally similar spin-flip-flop dynamic through coherent virtual phonon exchange. Thus, while it differs physically from exchange coupling, the effective result—entangling operations between spin qubits—is operationally analogous.
With typical parameters of $g_{\text{sp}}/2\pi\sim 5~{}\text{MHz}$ and detuning $\Delta/2\pi\sim 50~{}\text{MHz}$, the effective two-qubit gate rate lies in the range of 0.25–1 MHz, corresponding to entangling gate durations of approximately 0.5–1 $\mu\text{s}$. These timescales are well within the coherence times achievable in high-quality phononic cavities [43], enabling high-fidelity two-qubit operations in practical devices.
Phonon-mediated long-range interactions are fundamentally different from conventional exchange interactions, both in their physical mechanism and spatial characteristics. Exchange coupling relies on direct wavefunction overlap between neighboring qubits, necessitating sub-10 nm spacing and resulting in rapidly decaying, short-range interactions that limit architectural flexibility and scalability.
In contrast, phonon-mediated coupling exploits the virtual exchange of quantized lattice vibrations—phonons—confined within a PnCC. Each qubit interacts with a common phonon mode through strain-modulated spin–orbit coupling. Even when detuned from the phonon resonance, qubits can interact via virtual phonon exchange, enabling coupling over micrometer-scale distances without requiring physical proximity or wavefunction overlap. This architecture establishes a shared phononic “bus” capable of connecting multiple qubits across a chip.
A key advantage of this mechanism is its tunability. The interaction strength and phase can be dynamically modulated via qubit detuning, spin–orbit coupling parameters, and phononic cavity design. While both exchange and phonon-mediated interactions can generate entanglement, the latter more closely resembles dispersive interactions in cavity or circuit QED systems, rather than direct Coulomb- or tunneling-based exchange schemes.
@@
To support this form of coupling in Ge-based hole-spin qubits, we have designed a two-dimensional phononic crystal with an absolute bandgap centered around 4–6 GHz, with a bandwidth of 2–3 GHz. This frequency range corresponds to thermal energies near 0.3 K and is well-matched to dilution refrigerator environments operating at ∼$\sim 100\,\text{mK}$. At these temperatures, thermal occupation of acoustic modes can lead to decoherence. The PnCC structure acts as a spectral filter, suppressing thermally populated phonon modes outside the desired range while confining coherent phonon modes for interaction. This selective filtering enhances coherence and supports long-range, phonon-mediated qubit control in scalable Ge quantum architectures.
@@
In this work, we estimate the relative strength of spin–phonon coupling $g_{\text{sp}}$ by evaluating the spatial overlap between the phononic strain field and the qubit’s envelope wavefunction. To model the spatial envelope of the hole-spin qubit, we assume a Gaussian wavefunction of the form
$\psi(r)\propto\exp\left[-\frac{(r-r_{0})^{2}}{2\sigma^{2}}\right],$
where $r_{0}$ is the qubit center and $\sigma$ is the spatial width, related to the full width at half maximum (FWHM) by $\text{FWHM}=2\sqrt{2\ln 2}\,\sigma$. In our simulations, a FWHM of approximately 180 nm was used to reflect the lateral confinement of a gate-defined quantum dot.
To assess the sensitivity of spin–phonon coupling to wavefunction geometry, we varied $\sigma$ between 80 nm and 250 nm and found that the spin–phonon coupling strength $g_{\mathrm{sp}}$ exhibited a non-monotonic dependence, peaking when the wavefunction width closely matched the strain profile of the confined phonon mode. Spatial misalignment of the qubit center $r_{0}$ relative to the phonon mode maximum also significantly impacted the overlap integral, underscoring the importance of precise spatial co-localization between the PnCC mode and the qubit. These observations provide guidance for optimizing device geometry to maximize coupling efficiency.
Using this Gaussian approximation for the hole-spin qubit and the simulated strain profile of a 6 GHz confined phonon mode, we compute a normalized overlap integral of approximately $2.6\times 10^{-2}$. This value is over four orders of magnitude greater than that obtained at 20 GHz, reflecting the fact that lower-frequency phonons, which exhibit broader and smoother spatial profiles, are better matched to the spatial extent of gate-defined quantum dots. The improved spatial overlap reduces destructive interference within the qubit region and enables stronger, more coherent coupling between the spin and the phonon field.
@@ V Experimental Considerations
The experimental realization of phonon-coupled hole-spin qubits in Ge requires the coordinated integration of high-purity material growth, precision nanofabrication, and low-temperature quantum measurements. A critical enabler of this effort is the availability of detector-grade, high-purity Ge crystals grown in-house at USD, with impurity concentrations below $10^{10}~{}\text{cm}^{-3}$. This level of material purity significantly reduces charge noise and impurity scattering—key sources of qubit decoherence [13, 14, 15, 16, 17, 18, 19]—and establishes a realistic and practical foundation for future experimental demonstrations.
It is important to emphasize that this in-house capability is not intended to imply exclusivity. Rather, it highlights USD’s vertically integrated approach to crystal synthesis and purification. While this infrastructure enables us to tailor Ge substrates specifically for quantum applications, similar ultra-high-purity Ge crystals can also be obtained from a few specialized vendors worldwide, such as ORECT, Mirion (formerly Canberra), Unimore, and PHDS. These institutions possess the requisite zone-refining and crystal-growth expertise to achieve comparable impurity levels. In contrast, most commercially available Ge wafers exhibit impurity concentrations on the order of $10^{14}~{}\text{cm}^{-3}$, which is four orders of magnitude higher than what is necessary for coherent quantum operations in phonon-coupled Ge-based architectures.
@@
Quantum dots are electrostatically defined using overlapping gate architectures fabricated by high-resolution electron-beam lithography. A high-$\kappa$ dielectric layer (e.g., HfO2) deposited via atomic layer deposition isolates the gates from the Ge surface. The gate stack typically includes accumulation gates to form the quantum channel, barrier gates to define tunnel junctions, and plunger gates to tune the dot occupancy. Vertical electric fields generated by the gates modulate the hole g-factor through Rashba spin-orbit interaction, enabling EDSR without requiring oscillating magnetic fields [7, 8].
@@
The proposed qubit architecture is designed to operate at cryogenic temperatures, typically below 100 mK, where thermal phonon populations are sufficiently suppressed to allow coherent control and long spin relaxation times. At higher temperatures, the thermal occupation of phonon modes increases exponentially, leading to elevated rates of phonon absorption and enhanced spin relaxation. Specifically, the spin–phonon coupling strength $g_{sp}$ remains relatively stable with temperature; however, the resulting relaxation time $T_{1}\propto 1+n_{B}(\omega,T)$, where $n_{B}$ is the Bose-Einstein distribution, decreases sharply due to thermally activated processes. The performance of the PnCC is also temperature-dependent: although the acoustic bandgap structure remains intact, elevated temperatures introduce additional phonon modes that may leak into the cavity, degrading the quality factor and coherence protection. Based on simulations and literature data, the architecture retains high performance up to approximately 200–300 mK, beyond which $T_{1}$ and Q-factor degrade significantly. Therefore, stable operation is best maintained within dilution refrigeration environments.
@@ VI Device Design Concept
Figure 10: Schematic cross-section of the proposed Ge-based hole-spin qubit platform. A single hole is confined in a gate-defined quantum dot formed within a strained Ge quantum well, grown atop a relaxed Si1-xGex buffer layer. The structure includes a cap layer and gate electrodes that enable vertical electrostatic confinement and all-electrical spin control. The epitaxial strain enhances heavy-hole character and spin–orbit coupling, supporting fast and tunable qubit operations. This geometry serves as the foundation for integrating PnCCs to enable phonon-mediated long-range qubit interactions. Crystallographic directions are indicated to reflect the device’s orientation relative to quantum confinement and gate alignment.
@@
Figure 11: Schematic illustration of a Ge-based hole-spin qubit integrated with a PnCC. The qubit is electrostatically confined by gate electrodes at the interface between the cap layer and a strained Ge quantum well, grown on a relaxed Si1-xGex buffer. The epitaxial strain enhances spin–orbit coupling and coherence. A lateral PnCC is positioned adjacent to the qubit to enable coupling between the localized hole spin and engineered acoustic phonon modes. The cavity acts as a spectral filter and mediator for coherent spin manipulation and long-range qubit interactions. The phononic mode is spatially aligned with the quantum dot to maximize spin–phonon coupling efficiency.
@@ VII.2 Quantum Dot Behavior and Spin-Orbit Coupling
where $g_{\text{eff}}(E_{z})$ is the electric-field-dependent effective g-factor, $\mu_{B}$ is the Bohr magneton, and $B$ is the applied magnetic field. The spin–phonon coupling strength $g_{sp}$ is calculated using the overlap integral formalism introduced in Refs. [37, 46], as expressed in Eq. (4). This approach captures the modulation of spin states by phonon-induced strain fields through spin–orbit interactions. The formulation aligns with the spatial overlap framework presented in Eq. (4), providing a consistent and quantitative method for evaluating spin–phonon coupling in Ge quantum dots.
Building on a multiband $\mathbf{k}\cdot\mathbf{p}$ framework and self-consistent Poisson–Schrödinger solvers [20, 22], the simulation framework described above was employed to model the dependence of $g_{\text{eff}}$ and $g_{sp}$ on the vertical electric field. Figure 12 shows that increasing $E_{z}$ reduces $g_{\text{eff}}$ from 2.0 to 1.3 and simultaneously enhances $g_{sp}$ from 0.5 to over 6 MHz. This tunability is critical for fast EDSR, as it allows dynamic optimization of qubit control fidelity and gate speed. Quantitative simulation results are summarized in Table 6 in the Appendix, confirming the trends observed in the figure.
@@ VII.3 Phonon Dynamics: Relaxation and Confinement
We employed the previously described simulation framework to model spin relaxation dynamics and phonon confinement in the presence of engineered phononic crystal structures. The longitudinal spin relaxation time $T_{1}$ is governed by Fermi’s golden rule [10, 11], as expressed in Eq. (2), which relates the relaxation rate to the phonon density of states and the square of the spin–phonon matrix element.
In the low-temperature acoustic limit, the phonon density of states scales as $D(\omega)\sim\omega^{2}$, resulting in shorter $T_{1}$ times at higher phonon frequencies [10, 11]. The matrix element $M_{sp}$ in Eq. (2) represents the strength of spin–phonon coupling and is directly derived from the strain-modulated spin–orbit Hamiltonian $H_{\text{so}}(u(\vec{r}))$, through the spatial overlap integral defined earlier in Eq. (4). This formulation provides a consistent physical framework linking spin relaxation rates to the engineered phononic environment and the spatial structure of the quantum dot wavefunction.
The spin–phonon matrix element $M_{\text{sp}}$, which governs the qubit relaxation rate in Eq. (2), is evaluated using the deformation potential formalism [10, 11]. Specifically, $M_{\text{sp}}$ arises from the strain-induced modulation of spin–orbit coupling in the valence band and is given by:
$M_{\text{sp}}=\left\langle\psi_{\uparrow}(\vec{r})\middle|H_{\text{so}}(u(\vec{r}))\middle|\psi_{\downarrow}(\vec{r})\right\rangle,$
as defined in Eq. (4). Here, $H_{\text{so}}(u(\vec{r}))$ denotes the strain-dependent spin–orbit Hamiltonian, and $u(\vec{r})$ is the phonon displacement field computed from finite-element simulations of confined phonon modes [37, 46]. The qubit wavefunctions $\psi_{\uparrow,\downarrow}$ are modeled as Gaussian envelopes with tunable width and spatial location, reflecting the electrostatic confinement profile of the Ge quantum dot [7].
The magnitude of $M_{\text{sp}}$ depends on several key parameters, including the spatial overlap between the phonon-induced strain and the qubit wavefunction, the degree of heavy-hole/light-hole mixing, the applied electric field (which modulates Rashba spin–orbit interaction), and the orientation of the external magnetic field relative to the crystallographic axes [11, 8]. In our simulation framework, $M_{\text{sp}}$ is numerically computed by integrating the product of the simulated phononic strain profile and the modeled wavefunction envelope over the dot region. This methodology provides realistic, geometry-aware estimates of the transition matrix element that controls both $T_{1}$ relaxation and phonon-mediated qubit dynamics.
@@
where $\omega$ is the mode frequency and $\Delta\omega$ the linewidth obtained from eigenmode simulations.
As shown in Figure 13, maximum coherence and cavity performance occur near 6 GHz, with simulated $T_{1}\approx 1$ ms and $Q\approx 18000$. The numerical values corresponding to these trends are listed in Table 7 in the Appendix, providing quantitative validation of the simulation results shown in the figure.